304278:

The temperature in Bloomington on Sunday was 32.6°C. On Monday, the temperature changed by -8.25°C. What was the temperature on Monday?

Extrinix:

\(32.6-8.25=temperature~on~Monday\) easiest way I can give it to you.

Extrinix:

If you are trying to find \(^\circ F\) do this: \(^\circ F = ^\circ \hspace{-1.45px}C \times 1.8 + 32\)

Extrinix:

@extrinix wrote:
If you are trying to find \(^\circ F\) do this: \(^\circ F = ^\circ \hspace{-1.45px}C \times 1.8 + 32\)
forgot to add, do this with the answer from my response above that.